# The early universe is ACT-ing warm

### v1

@@ III Conclusions
Where does the P-ACT-LB-BK18 data leave inflationary model-building? At face value, a tiny $r$ and $n_{s}$ very close to unity tells us that the standard models of renormalizable, monomial potentials seem to fall out of favour of observations. However, the deeper revelation is that observations seem to strike at the heart of the problem of protecting the flatness of the potential from quantum corrections. One can always find one or two-parameter extensions of current models of cold inflation which fit the data as more precise mappings of the CMB come in; however, it is time we introspect to ask if such extensions have an embedding in any UV-theory. Put another way, can one accentuate well-motivated models with such ad hoc couplings which are neither derived from any underlying theory nor protected from the effect of Planck-suppressed operators? Even when one takes the modern view of EFT that non-renormalizable potentials are, by themselves, not problematic and simply point towards a UV-cutoff to which scale it can be trusted, it makes little sense to pick and choose operators (of the same derivative order) such that a model becomes viable. A principled path forward for inflation seems to require a new physical input which would help reconciling theoretically-consistent models with observations.
@@
In this paper, we argue that WI is the right paradigm to do so. It is only natural that the inflaton field interacts with other fields during inflation and the only further assumption WI makes is regarding the strength of the dissipative terms generated in the process. However, one then finds that standard models of WI can easily fit the data where their cold counterparts failed. One might question why is the specific potentials chosen in this paper not equally fine-tuned. Our main argument is that the exact choice of potential is not as important in WI as it is in the cold case. This is precisely because of the presence of the extra dissipative terms. We could just as easily have chosen some other potential and much of our discussion would still go through. Moreover, we do choose a form of the potential which is renormalizable and thereby can be assumed to be protected from Planckian corrections. As can be clearly seen from Fig. 1, dissipative corrections turn even these monomial potentials viable in WI. Notice that although the strong dissipative regime would be preferred from theoretical considerations, these models produce the right spectrum even for small $Q$. However, future observations of the tensor-to-scalar ratio, or the lack of it, can be used as a signature of the strong vs. weak dissipative regime for WI.
@@
Naturally, WI is also not free from any constraints. Firstly, the dissipative coefficients need to be computed from first-principles (and that is what has been done for the two models under consideration here). More needs to be done regarding the microphysics of such dissipative interactions and how they can be derived from a full thermal field theory perspective in an accelerating background. This can put further bounds on the parameter space of WI. Moreover, $Q$ is not a completely unconstrained parameter even from observations. It is natural to expect that going to arbitrarily large values of the dissipative coefficient would lead to an unacceptably large $f_{NL}$ for the warm shape in the scalar bispectrum (we have made sure that going upto $Q\sim\mathcal{O}(100)$ is consistent with current bounds on the bispectrum Bastero-Gil:2019gao ). We plan to find the upper bound on $Q$ coming from the bispectrum, which would also restrict how large $n_{s}$ can be in WI.
@@
For completeness, we also mention the possibility that inflation itself might not be the right way to think about the early-universe. A few alternatives to inflation Brandenberger:1988aj ; Steinhardt:2001st ; Agrawal:2020xek ; Brahma:2022dsd ; Boyle:2022lcq , which are grounded in a fundamental UV-theory, might be better suited to explain observations. Curiously, many of them also postulate thermal fluctuations Nayeri:2005ck ; Brahma:2021tkh as in WI, thereby pointing towards a larger application of dissipative corrections going beyond inflation.
Acknowledgements.
@@
AB is partially funded by STFC. SB is supported in part by the Higgs Fellowship and by the STFC Consolidated Grant “Particle Physics at the Higgs Centre”. R.O.R. and G.S.R. acknowledge financial support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) - Finance Code 001. R.O.R. also acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 5 Conclusions
Our results in general also point out to possible implications of the P-ACT-LB-BK18 data as far as inflationary model-building is concerned. At face value, a tiny $r$ and $n_{s}$ very close to unity tell us that the standard models of renormalizable, monomial potentials seem to fall out of favor of observations. However, the deeper revelation is that observations seem to strike at the heart of the problem of protecting the flatness of the potential from quantum corrections. One can always find one- or two-parameter extensions of current models of cold inflation which fit the data as more precise mappings of the CMB come in. Yet even accepting that, it has not been the same model that is relevant as the data changes, but rather the model has changed, selected from effectively an infinite space of models. Moreover, it is time we introspect to ask if such extensions have an embedding in any UV-theory. Put another way, can one accentuate well-motivated models with such ad hoc couplings that are neither derived from any underlying theory nor protected from the effect of Planck-suppressed operators? Even when one takes the modern view of EFT that non-renormalizable potentials are, by themselves, not problematic and simply point towards a UV-cutoff to which scale it can be trusted, it makes little sense to pick and choose between operators (of the same derivative order) such that a model becomes viable. This just turns model building into a speculative exercise rather than based on a theoretically reasoned foundation [56]. A principled path forward for inflation seems to require a new physical input which would help reconciling theoretically-consistent models with observations.
@@
In this paper, we argue that WI is the right paradigm to do so. It is only natural that the inflaton field interacts with other fields during inflation and the only further assumption WI makes is regarding the strength of the dissipative terms generated in the process. However, one then finds that standard models of WI can easily fit the data where their cold counterparts failed. One might question why the specific potential chosen in this paper is not equally fine-tuned. Our main argument is that the exact choice of potential is not as important in WI as it is in the cold case. This is precisely because of the presence of the extra dissipative terms. We could just as easily have chosen some other potential and much of our discussion would still go through. Moreover, we do choose a form of the potential which is renormalizable and thereby can be assumed to be protected from Planckian corrections. As can be clearly seen from Fig. 1, dissipative corrections make even these monomial potentials viable in WI. Note that although the strong dissipative regime would be preferred from theoretical considerations, these models produce the right spectrum even for small $Q$. Indeed, the results presented here simply confirm that strong dissipative regimes (of warm little models) are preferred by data and are not obtained by any fine-tuning to fit new observations. However, future observations of the tensor-to-scalar ratio, or the lack thereof, can be used as a signature of the strong vs. weak dissipative regime for WI.
@@
Naturally, WI is also not free from any constraints. Firstly, the dissipative coefficients need to be computed from first-principles and that is what has been done for WLI model considered here. More needs to be done regarding the microphysics of such dissipative interactions and how they can be derived from a full thermal field theory perspective in an accelerating background. (A possible application of this is the question of whether the inflaton thermalizes or not, something that we have considered, at least qualitatively in the appendix.) This can put further bounds on the parameter space of WI. Moreover, $Q$ is not a completely unconstrained parameter, even from observations. It is natural to expect that going to arbitrarily large values of the dissipative coefficient would lead to an unacceptably large $f_{NL}$ for the warm shape in the scalar bispectrum (we have made sure that going up to $Q\sim\mathcal{O}(100)$ is consistent with current bounds on the bispectrum [30]). We plan to find the upper bound on $Q$ coming from the bispectrum, which would also restrict how large $n_{s}$ can be in WI.
@@
For completeness, we also mention the possibility that inflation itself might not be the right way to think about the early universe. Some alternatives to inflation [74, 75, 76, 77, 78], which are grounded in a fundamental UV-theory, might be better suited to explain observations. Interestingly, many of them also postulate thermal fluctuations [79, 80] as in WI, thereby pointing toward a larger application of dissipative corrections that go beyond inflation.
Acknowledgments
@@
AB is partially funded by STFC. SB is supported in part by the Higgs Fellowship and by the STFC Consolidated Grant “Particle Physics at the Higgs Centre”. R.O.R. acknowledges financial support by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. The work of G.S.R. is supported by a PhD scholarship from FAPERJ. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Appendix A Thermalization of the inflaton in the WLI model
Here, we provide estimates to justify whether the inflaton could thermalize in the WLI model, that is, whether $n_{*}$ in eq. (2.6) can assume a Bose-Einstein distribution, $n_{*}\equiv n_{\rm BE}$. The answer for this question is, of course, strongly dependent on the details of the microphysics of the specific WI model under study. This would involve, for example, whether scattering rates of the inflaton with other fields could happen at a rate $\Gamma_{\rm scatt}$ fast enough compared to the expansion rate, $\Gamma_{\rm scatt}>H$. The condition $\Gamma_{\rm scatt}>H$ is usually assumed to justify thermalization of radiation particles in cosmology [81], although a full justification can only be appropriately given through either an appropriate lattice simulation of the full dynamics, or a dedicated kinetic Botzmann equation study for the occupation number for the inflaton in WI. A previous general analysis of the latter have first been given in ref. [45]. The results obtained there in fact indicated that thermalization, in the sense that the distribution function can assume a thermal equilibrium form, is justified when the rate $\Gamma$ is of order or greater than the expansion rate HH [^5] In fact, the results in ref. [45] have shown that even when $\Gamma<H$ and the system is out-of-equilibrium, the distribution function can assume a form proportional to the thermal one, with a prefactor given by $A\Gamma/(3H+\Gamma)$, where $A$ is a numerical coefficient satisfying $A\sim{\cal O}(1)$.. In the absence of a more complete analysis for the thermalization of the inflaton, it is common in most applications in WI to typically consider both possibilities — non thermalized inflaton perturbations, $n_{*}=0$, or fully thermalized, $n_{*}=n_{\rm BE}$ — in the expression for the scalar of curvature power in WI, eq. (2.5). When presenting all our results, we have assumed the same philosophy. However, we can take advantage of the fact that, since we know the details of the microphysical WLI model, we know all relevant interactions involving the inflaton field eq. (3.1) and hence can estimate the relevant scattering rates that contribute to the thermalization of the inflaton.
In our estimates below, we will focus on the direct interactions of the inflaton with the light fields $\chi_{1,2}$ and $\psi_{1.2}$. We will also consider, for simplicity, only the $2\to 2$ processes that come from these interactions. From eq. (3.1), we have that the direct interactions of the inflaton field with $\chi_{1,2}$ and $\psi_{1.2}$ are
$\mathcal{L}_{\rm int}$ $=$ $-g_{\phi\psi}M\cos(\phi/M)\bar{\psi}_{1L}\psi_{1R}-g_{\phi\psi}M\sin(\phi/M)\bar{\psi}_{2L}\psi_{2R}$ (A.1)
- $g_{\phi\chi}^{2}M^{2}\cos^{2}(\phi/M)|\chi_{1}|^{2}-g_{\phi\chi}^{2}M^{2}\sin^{2}(\phi/M)|\chi_{2}|^{2}.$
When shifting the inflaton field around its background value and perturbations, $\phi\to\phi+\delta\phi$, we find that the relevant vertices for the processes involving the inflaton perturbations up to second order in $\delta\phi$ are
$\mathcal{L}_{\rm int}(\delta\phi,\chi_{i},\psi_{i},\bar{\psi}_{i})$ $=$ $g_{\phi\psi}\sin(\phi/M)\delta\phi\bar{\psi}_{1L}\psi_{1R}-g_{\phi\psi}\cos(\phi/M)\delta\phi\bar{\psi}_{2L}\psi_{2R}$ (A.2)
$+$ $g_{\phi\chi}^{2}M\sin(2\phi/M)\delta\phi\left(|\chi_{1}|^{2}-|\chi_{2}|^{2}\right)$
$+$ $\frac{g_{\phi\psi}}{2M}\cos(\phi/M)\delta\phi^{2}\bar{\psi}_{1L}\psi_{1R}+\frac{g_{\phi\psi}}{2M}\sin(\phi/M)\delta\phi^{2}\bar{\psi}_{2L}\psi_{2R}$
$+$ $g_{\phi\chi}^{2}\cos(2\phi/M)\delta\phi^{2}\left(|\chi_{1}|^{2}-|\chi_{2}|^{2}\right).$
From the interaction Lagrangian density eq. (A.2) we read off the vertices:
${\rm Yukawa\;(linear\;in\;}\delta\phi):\qquad{\cal L}\supset y_{1}\,\delta\phi\,\bar{\psi}_{1L}\psi_{1R}-y_{2}\,\delta\phi\,\bar{\psi}_{2L}\psi_{2R},$
${\rm trilinear\;(}\delta\phi\chi\chi\!):\qquad{\cal L}\supset\kappa_{\chi}\,\delta\phi\,\left(|\chi_{1}|^{2}-|\chi_{2}|^{2}\right),$
${\rm contact\;(}\delta\phi^{2}\bar{\psi}\psi\!):\qquad{\cal L}\supset\lambda_{\psi,1}\,\delta\phi^{2}\bar{\psi}_{1L}\psi_{1R}+\lambda_{\psi,2}\,\delta\phi^{2}\bar{\psi}_{2L}\psi_{2R},$
${\rm contact\;(}\delta\phi^{2}\chi^{2}\!):\qquad{\cal L}\supset\Lambda_{\chi}\,\delta\phi^{2}\left(|\chi_{1}|^{2}-|\chi_{2}|^{2}\right),$ (A.3)
where we have defined the couplings
$y_{1}\equiv g_{\phi\psi}S,\;y_{2}\equiv g_{\phi\psi}C,$
$\kappa_{\chi}\equiv g_{\phi\chi}^{2}MS_{2},$
$\lambda_{\psi,1}\equiv\frac{g_{\phi\psi}}{2M}C,\;\lambda_{\psi,2}\equiv\frac{g_{\phi\psi}}{2M}S,$
$\Lambda_{\chi}\equiv g_{\phi\chi}^{2}C_{2},$ (A.4)
and where $C,\,S,\,C_{2}$ and $S_{2}$ are defined as $S\equiv\sin(\phi/M),\,C\equiv\cos(\phi/M),\,S_{2}\equiv\sin(2\phi/M),\,C_{2}\equiv\cos(2\phi/M)$, with $\phi$ denoting the background inflaton field.
Figure 6: Feynman diagrams for the tree-level $2\to 2$ scattering processes originating from the interaction terms in eq. (A.2). The top diagrams come from contact interactions, while the bottom diagrams come from the trilinear vertex involving the scalars $\delta\phi$ and $\chi_{i}$ and from the Yukawa vertex between $\delta\phi$ and $\psi_{i},\,\bar{\psi}_{i}$. A dashed line represents the $\delta\phi$ excitation, a dotted line represents the complex scalars $\chi_{i}$ and the solid lines the fermions $\psi_{i}$.
Examples of tree-level $2\to 2$ scattering processes originating from the interaction terms in eq. (A.3) are shown in fig. 6. Although a detailed computation of all thermal scattering rates that can contribute to the possible thermalization of the inflaton excitation $\delta\phi$ is beyond the scope of this paper, here we estimate the rates for some of the main processes leading to thermalization of the inflaton particle states.
Considering an arbitrary $2\rightarrow 2$ process involving relativistic particles, we have that the differential cross-section (for the $s$-channel) is defined as [82]:
${d\sigma\over d\Omega}={|\mathcal{M}|^{2}\over 64\pi^{2}s}~,$ (A.5)
where $\sqrt{s}$ is the centre-of-mass energy. We can make estimates of the magnitude of the different processes shown in fig. 6 by neglecting the angular correlations and we can take typical momenta and energies, both for external and internal legs, to be of the order of the temperature of the thermal bath, where we also assume that all particles are relativistic, with $m_{\phi},\,m_{\chi_{1,2}},\,m_{\psi_{1,2}}\ll T$. For the parameters considered in the text, these conditions are well satisfied. We can then write the cross-section of each one of the processes as
$\sigma_{i}\sim{|\mathcal{M}_{i}|^{2}\over 16\pi T^{2}}~,$ (A.6)
where $|\mathcal{M}_{i}|^{2}$ is the averaged amplitude square of the process. The rate is then given by $\Gamma_{i}=\langle\sigma_{i}n_{i}v\rangle$, where for relativistic particles we have $n_{i}=f_{i}\zeta(3)T^{3}/\pi^{2}$ (where $\zeta(3)\simeq 1.20206$), with $f_{i}$ denoting the number of degrees of freedom (e.g. $f_{\chi_{i}}=2$ for the complex scalars $\chi_{1,2}$, etc), and $v\sim 1$. Hence, at leading-order, we find the tree-level estimates valid in the relativistic regime $T\gg m_{\phi,\chi,\psi}$:
$\Gamma_{\psi_{i}\psi_{i}\leftrightarrows\delta\phi\delta\phi}$ $\simeq\sum_{i=1,2}\frac{\zeta(3)}{4\pi^{3}}\left(y_{i}^{4}T+2\lambda_{\psi_{i}}^{2}T^{3}\right)\simeq\frac{\zeta(3)}{16\pi^{3}}g_{\phi\psi}^{2}T\left(3g_{\phi\psi}^{2}+2\frac{T^{2}}{M^{2}}\right),$ (A.7)
$\Gamma_{\delta\phi\psi_{i}\leftrightarrows\delta\phi\psi_{i}}$ $\simeq\sum_{i=1,2}\frac{3\zeta(3)}{8\pi^{3}}\left(y_{i}^{4}T+\lambda_{\psi_{i}}^{2}T^{3}\right)\simeq\frac{\zeta(3)}{16\pi^{3}}g_{\phi\psi}^{2}T\left(3g_{\phi\psi}^{2}+\frac{T^{2}}{M^{2}}\right),$ (A.8)
$\Gamma_{\chi_{i}\chi_{i}\leftrightarrows\delta\phi\delta\phi}$ $\simeq\sum_{i=1,2}\frac{\zeta(3)}{4\pi^{3}}T\left(\Lambda_{\chi_{i}}^{2}+\frac{\kappa_{i}^{4}}{T^{4}}\right)\simeq\frac{\zeta(3)}{4\pi^{3}}g_{\phi\chi}^{4}T\left(1+g_{\phi\chi}^{4}\frac{M^{4}}{T^{4}}\right),$ (A.9)
$\Gamma_{\delta\phi\chi_{i}\leftrightarrows\delta\phi\chi_{i}}$ $\simeq\sum_{i=1,2}\frac{\zeta(3)}{4\pi^{3}}T\left(\Lambda_{\chi_{i}}^{2}+\frac{\kappa_{i}^{4}}{T^{4}}\right)\simeq\frac{\zeta(3)}{4\pi^{3}}g_{\phi\chi}^{4}T\left(1+g_{\phi\chi}^{4}\frac{M^{4}}{T^{4}}\right),$ (A.10)
where in the last step in the above equations, we have taken the average of the oscillatory terms using that $\phi\gg M$, which is valid for the model under consideration. For the parameters considered in section 4.2, our only free parameter is the inflaton-$\chi$ coupling $g_{\phi\chi}$, which can then be used to control the magnitude of the total rate $\Gamma_{\rm scatt}$, given by the sum of the rates (A.7)-(A.10).
Figure 7: The total rate for the tree-level $2\to 2$ processes divided by the Hubble rate in the case of $b=20$. The results are computed at the Hubble-exit point.
In fig. 7, we compare the magnitude of the total rate $\Gamma_{\rm scatt}$ with the Hubble rate. The results shown in fig. 7 are indicative that thermalization of the inflaton excitations can occur when the coupling $g_{\phi\chi}$ is sufficiently large and/or in the strong dissipative regime $Q\gg 1$. These results should also be seen as a lower-order estimate since we are only including lowest order tree-level processes and not including loop-mediated processes, which can be of same order, and also inelastic processes like $2\to 3$, $2\to 4$, etc. All of these processes will contribute to the thermalization of the inflaton excitations.